Project

General

Profile

Actions

MOM Report 2016 06 28

Expert start-up

list of issues found

  • Procedure to access the mezzanine if the cooling channel magnets are on: calculations redone by Josef. Update of the procedure.
  • Q1: power supply was not starting properly and has rebooted itself: intervention during the morning. Replaced the PS. Right polarity?
  • CDB: issue writing in the CDB the new tables: FIXED
  • Q4: the back of the PS was found open by the BLOC
  • ALH: wrong alarm limit for the water flow in Q6: FIXED
  • TOF2: current trip in TOF2_103: emailed Maurizio et al. One PMT has been left at a lower voltage (1700v).
  • Data mover issue after the re-installation of miceonrec01: SOLVED and TESTED

report

  • DC/MOM - MOM/DC handovers
  • Yordan has updated the DAQ software: Hall probes readout are now part of the binary data. Unpacker updated.
  • Online monitoring and Online reconstruction working well (tracker plots are noisy and await a new calibration)
  • Offline reconstruction is fine.
  • Target has been run.
  • No problems with RC (minor known issue)
  • MOM has been trained by JG for the helium bottle checks.
  • Beam dump done in the afternoon.
  • Reference run. Solved an issue with the EMR DAQ (restarted the crate).

Current operational conditions

  • Hall locked, no absorber

Open issues

  • TOF2 HV crate
  • Monitoring GUI for the tracker volumes needs to be added in the AppLauncher
  • SSH enable/disable mimic not implemented yet.

Updated by Franchini, Paolo over 7 years ago ยท 10 revisions